# Collapsing in polygonal dynamics

### v2

@@ 3.2 Corollaries for closed polygons
Since the scaling symmetry acts by birational transformations, then the coefficients of the infinitesimal monodromy are rational functions in $p_{i}^{(j)}$. Because of corollary 3.4, a potential collapse point $q=[q^{(0)}:\dots:q^{(d)}]$ is an eigenline of this $(d+1)\times(d+1)$ matrix. To find it, one needs to solve a linear system with $(d+1)$ equations involving the coefficient of the matrix, hence the result. ∎
In the case of the pentagram map and its generalizations, we can apply the corollaries. Note that convexity is not involved. This answers the questions raised by Schwartz in [8, §9.2], at least over $\mathbb{C}$.
Corollary 3.6.
Let $k$ be an algebraicly closed complete valued field. For the pentagram map and its generalisations in $\mathbb{P}^{2}(k)$, the collapse points are fixed points of Glick’s operator.
Proof.
The generalisation of Glick’s lemma to the cousins of the pentagram maps in $\mathbb{P}^{2}$ is achieved by [8, Lemma 9.6]. Aboud and Izosimov proved in [7] that Glick’s operator coincides with the infinitesimal monodromy, and argument in fact holds on any complete field. Combined with corollary 3.4, we get the result. ∎
Remark 4.
It would be interesting to consider a similar statement in the setting of polygonal dynamics on $\mathbb{R}^{d}$ with the group $\mathrm{Isom}^{+}(\mathbb{R}^{d})$. For instance in $\mathbb{R}^{2}$, Benoist and Hulin prove in [26] that when $n=4$, the iteration of some element of the folding group is generically conjugated, on the moduli space, to a translation on an elliptic curve. This matches definition 2.4 we gave for integrability. The authors conclude that, on the initial space, the polygons drifts towards a point on the line at infinity, following some “drifting vector”. When seen on the projective plane, the fixed points of positive isometries are on the line at infinity, which is coherent with our conjecture.
4 Polygonal dynamics in $\mathbb{P}^{1}$
For any polygonal dynamic in $\mathbb{P}^{1}$, we can make a precise version of corollary 3.4 and corollary 3.5. This is presented is theorem 4.2.
We will apply it to three kinds of polygonal dynamics: the leapfrog map, the “flat” cross-ratio and the newly defined “stair-case” cross-ratio dynamics (the only ones we know). The two last ones are linked to the theory of discrete holomorphic functions (see [12, §6]) and are endowed with a discrete curvature, which is an $n$-periodic sequence $\mu=(\mu_{i})_{i\in\mathbb{Z}}$, where each $\mu_{i}\in\mathbb{G}_{m}$ should be thought to live on the edge linking $p_{i}$ and $p_{i+1}$. They are left unchanged by the action of $\mathbb{P}\mathrm{GL}_{2}$, but can be modified by the scaling symmetry.
For any $n$-gon in $\mathbb{P}^{1}$, define its cross-ratio coordinates to be:
$c_{i}:=[p_{i},p_{i+1},p_{i-1},p_{i+2}]=\frac{p_{i}-p_{i-1}}{p_{i}-p_{i+2}}\frac{p_{i+1}-p_{i+2}}{p_{i+1}-p_{i-1}},\quad i\in\mathbb{Z}.$
They are $n$-periodic and invariant under projective transformations.
Remark 5.
Sometimes another convention is taken for the cross-ratio, which is:
$(a,b,c,d)=\frac{a-b}{b-c}\frac{c-d}{d-a}.$
We have the relation:
[a,b,c,d]=1(a,b,c,d).[a,b,c,d]=1-(a,b,c,d).
In order to stay coherent with [10], we’ll stick to the first convention.
Here’s a lemma that retrieves the monodromy of any polygon in $\mathbb{P}^{1}$. Its proof can be found in the original paper.
Lemma 4.1 – [10, Lemma 3.2].
Let $P=((p_{1},\dots,p_{n}),M)$ be a $n$-gon in $\mathbb{P}^{1}$ and $c_{i}=[p_{i},p_{i+1},p_{i-1},p_{i+2}]$ be its cross-ratio coordinates. Up to conjugation, suppose that $p_{0}=1,\penalty 10000\ p_{1}=\infty,\penalty 10000\ p_{2}=0$. Then in this projective chart, the monodromy matrix is:
$M=\left({\begin{array}[]{cc}0&c_{1}\\ -1&1\\ \end{array}}\right)\dots\left({\begin{array}[]{cc}0&c_{n}\\ -1&1\\ \end{array}}\right).$
When we have a scaling symmetry, this gives us a formula for the infinitesimal monodromy. Combined with corollary 3.4, this tells us what can be the collapse points for the dynamics.
@@ 4 Polygonal dynamics in $\mathbb{P}^{1}$
Suppose to have a polygonal dynamic on $\mathbb{P}^{1}$ with a scaling symmetry $\rho$. Let $P$ be a closed $n$-gon, and $c_{i}(z)$ the cross-ratio coordinates of $P_{z}$. Then the infinitesimal monodromy of $P$ is
$M^{\prime}=\begin{pmatrix}n/2+J&-K\\ I&n/2-J\end{pmatrix}$,
@@
$I=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{1}{p_{i}-p_{i+1}},\quad J=\frac{1}{2}\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{p_{i}+p_{i+1}}{p_{i}-p_{i+1}},\quad K=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{p_{i}p_{i+1}}{p_{i}-p_{i+1}}.$
@@
A similar matrix already appeared in [10, §6.2.1]. This answers the question asked by Arnold and Arreche at the end of [9], about the link between the pentagram map and cross-ratio dynamics.
@@
$M^{\prime}$ $=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{1}{p_{i}-p_{i+1}}\begin{pmatrix}p_{i}&-p_{i}p_{i+1}\\ 1&-p_{i+1}\end{pmatrix}$,
$\chi_{P}(X,Y)$ $=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{(X-p_{i}Y)(X-p_{i+1}Y)}{p_{i}-p_{i+1}}.$
Remark 8.
The dynamic of any closed 3-gon is trivial on the moduli space for any system in $\mathbb{P}^{1}$, since the action of $\mathbb{P}\mathrm{GL}_{2}$ is 3-transitive. So on the initial space, the iteration of $T$ is equivalent to the iteration of a projective transformation.
In some systems (like the flat cross-ratio dynamics [10, Lemma 5.3]), the iteration of $T^{2}$ on closed 4-gon is also trivial on the moduli space. This is reminiscent of classical results on the pentagram map (see [4, §2]). It provides one more evidence in favour of conjecture 3.2.
@@
We proceed as in the proof of [7, Prop. 4.1]. By applying the scaling symmetry, we deform $M$ into
$M_{z}=\left({\begin{array}[]{cc}0&c_{1}(z)\\ -1&1\\ \end{array}}\right)\dots\left({\begin{array}[]{cc}0&c_{n}(z)\\ -1&1\\ \end{array}}\right).$
For $1\leq i\leq n$, define $M_{z,i}$ to be the product of the first $i$ matrices (so $M_{z,n}=M_{z}$), and set $M_{z,0}=\mathrm{Id}$. Note that
$\left({\begin{array}[]{cc}0&c_{i}(z)\\ -1&1\\ \end{array}}\right)\dots\left({\begin{array}[]{cc}0&c_{n}(z)\\ -1&1\\ \end{array}}\right)=M_{z,i-1}^{-1}M_{z}.$
This gives that
$\frac{dM_{z}}{dz}=\left(\sum_{i=1}^{n}M_{z,i-1}\left({\begin{array}[]{cc}0&c_{i}^{\prime}(z)\\ 0&0\\ \end{array}}\right)M_{z,i}^{-1}\right)M_{z}.$
Since $M_{1}=\mathrm{Id}$, then the infinitesimal monodromy is
$M^{\prime}:=\left.\frac{dM_{z}}{dz}\right|_{z=1}=\sum_{i=1}^{n}S_{i},\quad\text{where }S_{i}:=M_{1,i-1}\left({\begin{array}[]{cc}0&c_{i}^{\prime}(1)\\ 0&0\\ \end{array}}\right)M_{1,i}^{-1}.$
Since we chose the projective chart where $p_{0}=1,p_{1}=\infty,p_{2}=0$, we obtain the lifts $V_{i}\in k^{2}\setminus\{(0,0)\}$ such that
$V_{0}=\begin{pmatrix}1\\ 1\end{pmatrix},V_{1}=\begin{pmatrix}1\\ 0\end{pmatrix},V_{2}=\begin{pmatrix}0\\ 1\end{pmatrix}$.
By proceeding like in [10, Lemma 3.2], we get that the matrices act this way:
$M_{1,1}:\begin{cases}V_{0}&\mapsto c_{1}V_{1}\\ V_{1}&\mapsto-V_{2}\\ V_{2}&\mapsto V_{3}\\ \begin{pmatrix}c_{2}\\ 1\end{pmatrix}&\mapsto V_{4}\\ \end{cases}.$
Indeed, the first three lines are given by direct computation, and the last one comes from the observation that:
$c_{2}=[1,\infty,0,c_{2}]=[p_{1},p_{2},p_{3},M_{1,1}(c_{2})].$
By definition of $c_{2}$ we obtain that $M_{1,1}(c_{2})=p_{4}$, which means that:
$M_{1,1}\begin{pmatrix}c_{2}\\ 1\end{pmatrix}=V_{4}.$
Similarly, we get by recurrence that for $2\leq i\leq n$:
$M_{1,i}:\begin{cases}V_{0}&\mapsto-c_{i}V_{i}\\ V_{1}&\mapsto-V_{i+1}\\ V_{2}&\mapsto V_{i+2}\\ \begin{pmatrix}c_{i+1}\\ 1\end{pmatrix}&\mapsto V_{i+3}\end{cases}.$
Since $p_{i}\neq p_{i+1}$, then $(V_{i},V_{i+1})$ forms a basis of $k^{2}$. By a simple computation we get:
$S_{i}:\begin{cases}V_{i}&\mapsto\frac{c_{i}^{\prime}(1)}{c_{i}}V_{i}\\ V_{i+1}&\mapsto\begin{pmatrix}0\\ 0\end{pmatrix}\end{cases}$.
We can decompose any $V\in k^{2}$ as
$V=\frac{|V,V_{i+1}|}{|V_{i},V_{i+1}|}V_{i}+\frac{|V_{i},V|}{|V_{i},V_{i+1}|}V_{i+1},$
where $|,|$ denotes the determinant of the $2\times 2$ matrix made out the column vectors. In this basis, the linear map associated to $S_{i}$ is:
$V\mapsto\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{|V,V_{i+1}|}{|V_{i},V_{i+1}|}V_{i}.$
Notice that this linear map commutes with the diagonal action of $\mathrm{GL}_{d+1}$ on the vectors. Indeed, the ratio of the determinants remains unchanged. Moreover, the previous linear map is independent of the choice of lifts $V_{i}$’s. So up to a projective transformation, we can suppose that $p_{i}\neq\infty$ for all $i$ and choose the lifts:
$V_{i}=\begin{pmatrix}p_{i}\\ 1\end{pmatrix}\quad\forall i\in\mathbb{Z}.$
In this choice of affine chart, the matrix expression of the linear map is:
$S_{i}=\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{1}{p_{i}-p_{i+1}}\begin{pmatrix}p_{i}&-p_{i}p_{i+1}\\ 1&-p_{i+1}\end{pmatrix}$.
By summing them, we conclude that:
$M^{\prime}=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{1}{p_{i}-p_{i+1}}\begin{pmatrix}p_{i}&-p_{i}p_{i+1}\\ 1&-p_{i+1}\end{pmatrix}$.
A small computation makes the quantities $I,J,K$ appear. Note that the eigenvectors of the matrix
$\begin{pmatrix}n/2+J&-K\\ I&n/2-J\end{pmatrix}$
are the same that the ones of
@@
The eigenvalues change, but it doesn’t matter for our purpose. Let $\begin{pmatrix}X\\ Y\end{pmatrix}$ be an eigenvector with eigenvalue $\lambda$, then:
@@
The quantities $I,J,K$ are invariant by the dynamic. Furthermore, the quantity $\Delta=:J^{2}-IK$ is invariant under the action of $\mathbb{P}\mathrm{GL}_{2}$, but the quantities $I,J,K$ are not. They change under the action of the generators as the following:
@@
$=[I,-\chi^{\prime}_{P}(\lambda,1),\chi_{P}(\lambda,1)],\qquad\penalty 10000\ \lambda\in k,$
$(z\mapsto\lambda z)\cdot[I,2J,K]$ $=[I\lambda^{-1},2J,K\lambda],\qquad\qquad\qquad\lambda\in k^{*},$
@@
The invariance of $I,J,K$ under the dynamic simply comes from the invariance of the infinitesimal monodromy.
The quantity $\Delta=J^{2}-IK$ is the determinant of the polynomial $\chi_{P}$, and so it is projectively invariant.
The action of projective transformations does not change the cross-ratio coordinates, but only the $p_{i}$’s. The modifications of $I,J,K$ under the action of the generators are given by a simple computation. ∎
Some objects keep reappearing in each polygonal dynamic in $\mathbb{P}^{1}$, but we couldn’t make a unifying statement. Let’s conclude this section with open questions.
@@
Define the following 1-form and (usually presymplectic) 2-form:
$\lambda=\frac{1}{2}\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{dp_{i}+dp_{i+1}}{p_{i+1}-p_{i}},\quad\Omega=d\lambda=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{dp_{i}\wedge dp_{i+1}}{(p_{i}-p_{i+1})^{2}}.$
Long specific computations show that in each system, $\Omega$ is invariant through the dynamic, meaning $T^{*}\Omega=\Omega$ (see [10, Thm. 15] and [11, Cor. 5.15]). Could we prove it for any polygonal dynamic in $\mathbb{P}^{1}$?
Moreover, if we consider the generators of the infinitesimal action of $\mathbb{P}\mathrm{GL}_{2}$
$u$ $=\sum_{k\in\mathbb{Z}/n\mathbb{Z}}\frac{\partial}{\partial p_{k}},$ $v$ $=\sum_{k\in\mathbb{Z}/n\mathbb{Z}}p_{k}\frac{\partial}{\partial p_{k}},$ $w=$ $\sum_{k\in\mathbb{Z}/n\mathbb{Z}}p_{k}^{2}\frac{\partial}{\partial p_{k}},$
then a little computation similar to the one done in [10, §6.2] shows that
$i_{u}\Omega$ $=dI,$ $i_{v}\Omega$ $=dJ,$ $i_{w}\Omega$ $=dK+\sum_{i=1}^{n}\frac{1}{2}(\frac{c_{i}^{\prime}(1)}{c_{i}}-\frac{c_{i-1}^{\prime}(1)}{c_{i-1}})dp_{i}.$
If all the $\frac{c_{i}^{\prime}(1)}{c_{i}}$ are equal [^10] For the systems considered here, it only happens for the flat cross-ratio dynamics with constant discrete curvature., then the last sum vanish and this tells us that $I,J,K$ are the hamiltonians of respectively $u,v,w$.
@@
For $1\leq k\leq\lfloor n/2\rfloor$, consider the quantities
$G_{k}(P)=\sum_{i_{1}<\dots<i_{k}}\frac{c_{i_{1}}^{\prime}(1)}{c_{i_{1}}}\dots\frac{c_{i_{k}}^{\prime}(1)}{c_{i_{k}}}\frac{(p_{i_{1}}-p_{i_{k}+1})(p_{i_{2}}-p_{i_{1}+1})\dots(p_{i_{k}}-p_{i_{k-1}+1})}{(p_{i_{1}}-p_{i_{1}+1})(p_{i_{2}}-p_{i_{2}+1})\dots(p_{i_{k}}-p_{i_{k}+1})}.$
They are projective invariant, since they are sums of multiratios. They are introduced for the flat cross-ratio dynamics in [10, §5.3], where the authors prove the invariance through the dynamic. They also link them to the concept of $\lambda$-length, and to logarithms of alternated perimeters coming from ideal polygons.
A long and painful computation (not reproduced here) shows that these quantities are also invariant for the staircase cross-ratio dynamic. But for the general case of polygonal dynamic over $\mathbb{P}^{1}$, we didn’t manage to prove the invariance. Maybe the answer could come from the $\lambda$-length interpretation, more detailed in [29], which would be coherent from the observation that
$\frac{c_{i}^{\prime}(1)}{c_{i}}=\left.\frac{d\mathrm{ln}(c_{i}(z))}{dz}\right|_{z=1}.$

### v3

@@ 4.1 The infinitesimal monodromy and the limit point
Since $\mathbb{P}\mathrm{GL}_{2}$ is $3$-transitive, then $\mathcal{P}_{3}$ is a point and $T$ has to act trivially.
Let $P$ be a closed $4$-gon. Because of [arnoldCrossratioDynamicsIdeal2022, Thm. 1], the trace of its monodromy in the projective chart $p_{0},p_{1},p_{2}=1,\infty,0$ is
$\mathrm{Tr}(M)=1-(c_{1}+c_{2}+c_{3}+c_{4})+c_{1}c_{3}+c_{2}c_{4}.$
But we have
$c_{2}$ $=[p_{2},p_{3},p_{1},p_{0}]=(1-c_{1}),$
$c_{3}$ $=[p_{3},p_{0},p_{2},p_{1}]=c_{1},$
$c_{4}$ $=[p_{0},p_{1},p_{3},p_{2}]=(1-c_{1}),$
so
$\mathrm{Tr}(M)=2c_{1}(c_{1}-1).$
The first cross-ratio of the iterates $T^{N}(P)$ always have to be one of the two roots of the polynomial
$X^{2}-X-\frac{\mathrm{Tr}(M)}{2},$
which forces $T^{2}=\mathrm{Id}$ on $\mathcal{P}_{4}$. ∎
We have an explicit formula of the infinitesimal monodromy (from theorem 2.10) and the degree $2$ polynomial equation verified by the limit point (from corollary 3.6).
@@
Let $T$ be a polygonal transformation in $\mathbb{P}^{1}$ admitting a scaling symmetry. Let $P=(p_{1},\dots,p_{n})$ be a closed $n$-gon and $(c_{i})$ be its cross-ratio coordinates. Then, its infinitesimal monodromy is
$M^{\prime}(1)=\frac{n}{2}\mathrm{Id}+\begin{pmatrix}J&-K\\ I&-J\end{pmatrix}$,
@@
$I=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}(1)}\frac{1}{p_{i}-p_{i+1}},\quad J=\frac{1}{2}\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}(1)}\frac{p_{i}+p_{i+1}}{p_{i}-p_{i+1}},\quad K=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}(1)}\frac{p_{i}p_{i+1}}{p_{i}-p_{i+1}}.$
@@
A similar matrix already appeared in [arnoldCrossratioDynamicsIdeal2022, §6.2.1]. This answers the question asked by Arnold and Arreche at the end of [arnoldSymmediansHyperbolicBarycenters2024], about the link between the pentagram map and cross-ratio dynamics.
@@
$M^{\prime}(1)$ $=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}(1)}\frac{1}{p_{i}-p_{i+1}}\begin{pmatrix}p_{i}&-p_{i}p_{i+1}\\ 1&-p_{i+1}\end{pmatrix}$,
$\chi_{P}(X,Y)$ $=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}(1)}\frac{(X-p_{i}Y)(X-p_{i+1}Y)}{p_{i}-p_{i+1}}.$
@@
Theorem 2.10 tells us the general form of the infinitesimal monodromy. As we said in the beginning of the section, we always normalize $c_{i,2}(t)=1$ (which forces $c_{i,2}^{\prime}(1)=0$) and identify $c_{i}(t)$ with $c_{i,1}(t)$ for every $t\in\mathbb{G}_{m}$. Thus we have
$M^{\prime}(1):p\mapsto\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}(1)}\frac{\mathrm{det}(p,p_{i+1})}{\mathrm{det}(p_{i},p_{i+1})}p_{i}.$
We can rewrite it as the matrix
$M^{\prime}(1)=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}(1)}\frac{1}{p_{i}-p_{i+1}}\begin{pmatrix}p_{i}&-p_{i}p_{i+1}\\ 1&-p_{i+1}\end{pmatrix}$,
which we recognize to be
$\frac{n}{2}\mathrm{Id}+\begin{pmatrix}J&-K\\ I&-J\end{pmatrix}$.
Because of remark 7, this represents the same matrix in $\mathfrak{pgl}_{2}$ than
@@
In both cases, an eigenvector $\begin{pmatrix}X\\ Y\end{pmatrix}$ with eigenvalue $\lambda$ verifies:
@@
4.2 Preserved quantities
We retrieve many preserved quantities. The first ones is a consequence of theorem 4.2.
@@ 4.2 Preserved quantities
The quantities $I,J,K$ are preserved by the dynamic, but not by projective transformations. They change under the action of the generators of $\mathbb{P}\mathrm{GL}_{2}$ in the following way:
@@
$=[I,-\chi^{\prime}_{P}(\lambda,1),\chi_{P}(\lambda,1)],\qquad\penalty 10000\ \lambda\in F,$
$(z\mapsto\lambda z)\cdot[I,2J,K]$ $=[I\lambda^{-1},2J,K\lambda],\qquad\qquad\qquad\lambda\in F^{*},$
@@
The invariance of $I,J,K$ under the dynamic simply comes from the invariance of the infinitesimal monodromy. The action of projective transformations does not change the cross-ratio coordinates, but only the $p_{i}$’s. The change of $I,J,K$ under the action of the generators are given by a simple computation.
We have
$\Delta=\begin{vmatrix}J&K\\ I&J\end{vmatrix}$,
so it is projective invariant since the determinant is unchanged by conjugation. ∎
If the scaling symmetry is linear, we have some good candidates for Hamiltonians. Following [arnoldCrossratioDynamicsIdeal2022, §3.1], a subset $I\subset\{1,\dots,n\}$ is said to be cyclically sparse if it doesn’t contain a pair of consecutive indices (understood modulo $n$). We abbreviate it $I\circlearrowleft$ sparse”.
Proposition 4.4.
Suppose that the scaling symmetry is linear. Let $P$ be a (twisted or closed) $n$-gon in $\mathbb{P}^{1}$. For $0\leq k\leq\lfloor n/2\rfloor$, define
$H_{k}=\sum_{\underset{|I|\geq k}{I\circlearrowleft\text{sparse}}}(-1)^{|I|}c_{I}\sum_{\underset{|J|=k}{J\subset I}}\frac{c_{J}^{\prime}}{c_{J}}.$
These quantities are preserved by the dynamic and are projective invariant. Moreover, we have for every $t\in\mathbb{G}_{m}$:
$\mathrm{Tr}(M(t))=\sum_{k=0}^{\left\lfloor n/2\right\rfloor}H_{k}(t-1)^{k}.$
Proof.
We know from [arnoldCrossratioDynamicsIdeal2022, Thm. 1] that for any polygon, the trace of its monodromy is
$\mathrm{Tr}(M(t))=\sum_{k=0}^{\left\lfloor n/2\right\rfloor}(-1)^{k}\sum_{\underset{|I|=k}{I\circlearrowleft\text{sparse}}}c_{I}(t).$
We obtain the desired formula by applying lemma 2.12. The $H_{k}$’s are invariant under the dynamic because the monodromy is, and under projective transformations because the trace is. ∎
We also retrieve many projective invariant quantities, which are a generalization of the $G_{k}$’s defined in [arnoldCrossratioDynamicsIdeal2022, §5.2]. It would be interesting to have an interpretation of them as alternating semi-perimeters, like in [arnoldCrossratioDynamicsIdeal2022, §5.3].
Proposition 4.5.
Suppose that the scaling symmetry is linear. Let $P$ be a closed $n$-gon in $\mathbb{P}^{1}$. For $1\leq k\leq\lfloor n/2\rfloor$, define
$G_{k}=\sum_{i_{1}<\dots<i_{k}}\frac{c_{i_{1}}^{\prime}(1)}{c_{i_{1}(1)}}\dots\frac{c_{i_{k}}^{\prime}(1)}{c_{i_{k}}(1)}\frac{(p_{i_{1}}-p_{i_{k}+1})(p_{i_{2}}-p_{i_{1}+1})\dots(p_{i_{k}}-p_{i_{k-1}+1})}{(p_{i_{1}}-p_{i_{1}+1})(p_{i_{2}}-p_{i_{2}+1})\dots(p_{i_{k}}-p_{i_{k}+1})}$
and $G_{0}=2$. These quantities are preserved by the dynamic and are projective invariant. Moreover, we have for every $t\in\mathbb{G}_{m}$:
$\mathrm{Tr}(M(t))=\sum_{k=0}^{\left\lfloor n/2\right\rfloor}\frac{G_{k}}{k!}(t-1)^{k}.$
Combining the two properties tells us that for every $0\leq k\leq\left\lfloor n/2\right\rfloor$, the quanties are related by
$H_{k}=\frac{G_{k}}{k!}.$
Proof.
By proposition 2.14, we have that for any $k\in\mathbb{N}$:
$M^{(k)}(1)$ $=\sum_{1\leq i_{1}<\dots<i_{k}\leq n}S_{i_{1}}\dots S_{i_{k}}$
$=\sum_{1\leq i_{1}<\dots<i_{k}\leq n}\prod_{l=1}^{k}\frac{c_{i_{l}}^{\prime}(1)}{c_{i_{l}}(1)}\frac{1}{p_{i_{l}}-p_{i_{l}+1}}\begin{pmatrix}p_{i_{l}}&-p_{i_{l}}p_{i_{l}+1}\\ 1&-p_{i_{l}+1}\end{pmatrix}$.
But by the same computation as in the proof of [arnoldCrossratioDynamicsIdeal2022, Thm. 10], we obtain that $M^{(k)}(1)=0$ if $k>\left\lfloor n/2\right\rfloor$, and otherwise
$\mathrm{Tr}(M^{(k)}(1))=G_{k}.$
We conclude with Taylor’s theorem. Their invariance is inherited from the one of the $H_{k}$’s. ∎
Let’s finish this section with some questions.
@@
Is there a automatic way to produce an invariant Poisson bracket for which the $H_{k}$’s are the Hamiltonians, which could prove Liouville integrability as in [arnoldCrossratioDynamicsIdeal2022] ?
@@
Define the following 1-form and (usually presymplectic) 2-form on the space of closed $n$-gons $\tilde{\mathcal{P}}_{n}$:
$\lambda=\frac{1}{2}\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}(1)}\frac{dp_{i}+dp_{i+1}}{p_{i+1}-p_{i}},\quad\Omega=d\lambda=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}(1)}\frac{dp_{i}\wedge dp_{i+1}}{(p_{i}-p_{i+1})^{2}}.$
Consider the generators of the infinitesimal action of $\mathbb{P}\mathrm{GL}_{2}$
$u$ $=\sum_{1\leq k\leq n}\frac{\partial}{\partial p_{k}},$ $v$ $=\sum_{1\leq k\leq n}p_{k}\frac{\partial}{\partial p_{k}},$ $w=$ $\sum_{1\leq k\leq n}p_{k}^{2}\frac{\partial}{\partial p_{k}}.$
A computation similar to the one done in [arnoldCrossratioDynamicsIdeal2022, §6.2] shows that
$i_{u}\Omega$ $=dI,$ $i_{v}\Omega$ $=dJ,$ $i_{w}\Omega$ $=dK+\sum_{i=1}^{n}\frac{1}{2}(\frac{c_{i}^{\prime}(1)}{c_{i}(1)}-\frac{c_{i-1}^{\prime}(1)}{c_{i-1}(1)})dp_{i}.$
If all the $\frac{c_{i}^{\prime}(1)}{c_{i}(1)}$ are equal, [^12] For the systems considered here, it only happens for the flat cross-ratio dynamics with constant discrete curvature. then the last sum vanishes and this tells us that $I,J,K$ are the hamiltonians of respectively $u,v,w$.
Long specific computations show that in each system, $\Omega$ is invariant through the dynamic, meaning $T^{*}\Omega=\Omega$ (see [arnoldCrossratioDynamicsIdeal2022, Thm. 15] and [gekhtmanIntegrableClusterDynamics2016, Cor. 5.15]). This fact seems to be true for any polygonal dynamic in $\mathbb{P}^{1}$. An idea to prove it would be to rephrase the dynamic as a cluster transformation and generalize the interpretation of $\lambda$-length made in [arnoldCrossratioDynamicsIdeal2022, §5.3] with our cross-ratio coordinates. [^13] This is coherent with the fact that $\frac{c_{i}^{\prime}(1)}{c_{i}(1)}=\left.\frac{d\mathrm{ln}(c_{i}(t))}{dt}\right|_{t=1}.$ Then, $\Omega$ could be the Weil–Petersson Kähler two-form presented in [pennerDecoratedTeichmullerTheory2012, §2.3].